# ManipVQA: Injecting Robotic Affordance and Physically Grounded Information into Multi-Modal Large Language Models

### v1

@@ III-C Task Formulation
III-D MLLM Finetuning Strategy
Model Architecture. We adopt the MLLM, SPHINX [13] as our primary architecture, with LLaMA2 [12] serving as the language backbone. Given the necessity for both global and local visual grounding in robotic tasks, we integrate the visual encoder from CLIP [17] to extract local semantic features and the Q-Former [31] for summarizing visual features. Spatial alignment is facilitated using projection layers, and global features are merged with local ones through channel-wise concatenation. We acknowledge that the standard image resolution for pre-trained visual encoders, typically $224\times 224$, is insufficient for detailed visual perception. This limitation is particularly significant for robotic affordance reasoning, which often requires fine-grained visual grounding of object parts, such as tool handles or machine buttons. To augment ManipVQA’s region-level grounding capabilities, we employ a sub-images patching strategy following  [13]. Specifically, we partition a $448\times 448$ image into four $224\times 224$ sub-images taken from each corner, thereby preserving intricate visual details. The resulting image tokens are then positioned before the language instructions to provide visual context for the ensuing prompts.
Finetuning Strategy. As elucidated in Sec. III-A, we model both affordance and physical concepts within natural language representations and training samples are formatted in line with the general VQA framework, as delineated in Sec. III-C. As a result, the training objective employs a unified cross-entropy loss, diverging from the approaches in  [6, 4]. To maintain the model’s broad visual reasoning proficiency, we amalgamate general visual reasoning exercises with tasks specific to robotics.
@@ IV-A Implementation Details
Training Details. We fine-tuned the ManipVQA model using the SPHINX framework [13] on eight NVIDIA A100 (80GB) GPUs. The fine-tuning was completed in a single epoch, which took approximately 6 hours. During this phase, the visual encoders were kept frozen to maintain the integrity of the pre-trained features. The pre-trained model was the SPHINX-1K, obtained from the official repository. Training was conducted with a batch size of 4 and a learning rate set to $2\times 10^{-5}$.
@@
TABLE III: Evaluation results on HANDAL. Task abbreviations are as follows: ”Obj.” for Complete Object Detection, ”Aff.” for Robotic Affordance Detection, and ”Gr” for Grounded Detection. The letter ”B” denotes bounding box format, and ”M” denotes mask representation format. Object abbreviations are listed in sequence: Hammer, Pliers-Fixed Joint, Pliers-Slip Joint, Pliers-Locking, Power Drill, Ratchet, Screwdriver, Wrench-Adjustable, Wrench-Combinational, Label, Measuring Cup, Mug, Pan, Spatula, Strainer, Utensil, and Whisk.
Connected With Robotic Policy. The objective of ManipVQA is to augment the generalizability of robotic control policies. It could be used in a language-free format during high-level decision-making for physically grounded knowledge. For affordance localization, while the initial bounding box identifies areas of potential manipulation, it may also encompass extraneous elements like background features. To achieve a more precise affordance map, we leverage the SAM-HQ variant [32], which uses the initial bounding box as a ”box prompt” for more accurate segmentation. Heuristic methods are then employed to determine the contact point. In our experiments, the ground truth (GT) surface normals, essential for rotation estimation, were presumed to be accessible either through RGB-D sensing or an alternative pre-trained model.
@@ IV-B Experimental Setup
Robotic Affordance Detection. Our evaluation is primarily conducted on the HANDAL dataset [10], assessing both bounding box average precision (AP) and pixel-wise segmentation AP. Unlike the baseline model in the HANDAL, which detects only the whole object, our ManipVQA is capable of identifying both the entire object and its manipulable parts, namely the affordance regions. Additionally, we compare our approach to LISA [33], which integrates LLM and SAM decoder.
Physical Concept Grounding. For evaluating physical concept grounding capabilities, we utilize the PhysObjects dataset [11]. We benchmark ManipVQA against PG-InstructBLIP [11], a fine-tuned version of InstructBLIP on PhysObjects, and the latest and most advanced MLLM, GPT-4V. Due to the limited localization capacity of GPT-4V, we pair it with Set-of-Mark [34], where bounding boxes and index numbers are explicitly annotated on the target objects within the input images.
General Affordance Grounding. Although ManipVQA is trained solely on a robotic affordance dataset, we are interested in its generalization capabilities on broader affordance grounding datasets, such as on AGD20K [27]. This exploration is motivated by the robust reasoning and generalization potential of LLMs. Our method is evaluated on AGD20K and follows its metrics, including KLD, SIM, and NSS. We compared with AffordanceLLM [6], Cross-View-AG [27], LOCATE [35] and 3DOI [36].
Robotic Manipulation Tasks. We further integrate ManipVQA with a basic robotic control policy, as detailed in Section IV-A, and use the manipulation success rate as a metric to gauge its practicality in robotic manipulation tasks using PartNet-Mobility within the SAPIEN [37]. Our experimental setup is modeled after ManipLLM [4], utilizing identical metrics. However, we employ ground truth (GT) surface normals from the simulator for rotation estimation. We compare our method with Where2Act [38], FlowBot3D [39], and ManipLLM [4]. For additional details on the experimental settings, please refer to ManipLLM [4].
@@ IV-C Results
Robotic Affordance Detection Evaluation. As illustrated in Table III, our ManipVQA achieves remarkable performance in the detection of both complete objects and their affordances in a unified framework. Enhanced by the SAM-HQ, our model attains superior results in tasks involving affordance detection and grounding with a mask. However, it falls short in complete object segmentation, which we attribute to a tendency of SAM-HQ to cause over-segmentation.
Figure 3: Illustration for the affordance detection and grounding tasks on the HANDAL Dataset [10]. The first four columns depict results on affordance detection, while the final two columns focus on the affordance grounding task. In the affordance grounding task, rather than providing explicit part names, a task description is given, specifying the intended use of the tool object. The blue masks are obtained with SAM-HQ while the red masks are the LISA’s output.
Physical Concept Grounding Evaluation. Table IV presents the evaluation results on the PhysObjects [11]. It is noteworthy that even the advanced multi-modal large language model GPT-4v encounters challenges with tasks that require an understanding of physical concepts. This struggle probably stems from its limited capacity for precise localization and a deficiency in visual physical reasoning. With the integration of SoM [34], the GPT-4v has better localization ability while its performance remains suboptimal. Our ManipVQA outperforms PG-InstructCLIP [11] which is also fine-tuned on PhysObjects, and we hypothesize that this enhanced performance can be attributed to the more powerful large language model and the ensemble of vision encoders deployed within our MLLM.
@@
General Affordance Grounding. Table V presents the evaluation results on the AGD20K dataset [27] using the Hard split as defined by [6]. Remarkably, although our ManipVQA framework is trained solely on robotic affordances, such as grasping, it still demonstrates competitive performance on the general affordance grounding task. Notably, our method achieves the highest NSS score but the lowest KLD, which we attribute to the tendencies of the SAM-series models to over-segment images. Additionally, the discrepancy between the ground truth represented as a heatmap, and our output, a segmentation mask, likely contributes to the lower KLD score. Our method also shows a promising ability to distinguish between fine-grained affordances associated with the same object class. For example, it can differentiate between Cut-with and Hold actions for a knife, which involve the blade and handle, respectively, as illustrated in Fig. 5.
Figure 4: Ilustration for the general affordance grounding tasks on the AGD20K [27]. Notably, the action Sit-On is unseen in our affordance training split. The green bboxes are our model’s output while the blue masks are obtained with SAM-HQ.
Method SIM normal-$\uparrow$ NSS normal-$\uparrow$ KLD normal-$\downarrow$
@@
TABLE V: General affordance evaluation results on AGD20K [27]. Notably, our ManipVQA has been trained exclusively on basic robotic affordances, such as ”grasping a handle”, whereas the AGD20K annotations encompass a more diverse range of human-centric affordances.
Figure 5: Robust reasoning ability of the system to differentiate fine-grained features within the affordances associated with the same object class.
Robotic Manipulation In Simulator. Table VI illustrates the zero-shot performance of our model in SAPIEN [37] when combined with a basic heuristic-based control policy. The model’s success is largely due to the retention of commonsense reasoning capabilities in the MLLM and the incorporation of affordance knowledge, which enables effective robotic manipulation without prior fine-tuning on the task-specific data.
Figure 6: Visualizatios of ManipVQA within the SAPIEN simulator. The first sub-image illustrates the experimental setup, wherein the model is tasked to predict the gripper’s contact point to enable the desired movement of a specific object part. The green bboxes represent our model’s predictions, while the blue masks acquired using SAM-HQ, indicating the target area. The contact point corresponds to the geometric center of this mask.
@@
TABLE VI: Performance evaluation within the SAPIEN simulator using PartNet-Mobility Dataset. Notably, while the baseline methods in this evaluation use distinct training and testing splits, our model achieves robust performance without fine-tuning on the SAPIEN samples.
@@ IV-D Further Analysis
Ablation Studies. To dissect the contributions of the designs in ManipVQA, we performed a series of ablation studies. The results are presented in Table VII. It was observed that the SOTA MLLM, SPHINX, cannot execute vision-based reasoning tasks in robotics without the ManipVQA dataset. Furthermore, without the mixture of general vision data during the fine-tuning process, there is a noticeable decline in its physical understanding and affordance reasoning capabilities. This suggests that the current ManipVQA dataset alone may not provide a sufficiently large or diverse sample for effectively fine-tuning an MLLM. Additionally, the absence of visual ensembles leads to a significant drop in the model’s ability to reason about affordances, likely because robotic affordance reasoning often requires detailed part-level understanding.
Manip.Data Vis. Ens Mix. Train Phy. Aff. Box Aff. Mask
@@
TABLE VII: Ablation Studies. ”Manip.Data” indicates the use of the ManipVQA dataset, ”Vis. Ens” represents the employment of a visual encoder ensemble, and ”Mix.Train” refers to the inclusion of a mixed general visual dataset during fine-tuning. ”Phy.” assesses the model’s physical grounding; and ”Aff.” denotes the model’s affordance reasoning capabilities.
Impact on Pre-existing Vision Reasoning Ability. When fine-tuning a model on a specialized dataset like ManipVQA, it is essential to consider the potential impact on its pre-established general vision reasoning skills. To investigate this, we evaluated the model on the val split of refCoco+ [28]. The post-fine-tuning accuracy was recorded at 81.8%. This performance is to be compared against the accuracy of the pre-trained model from  [2] at 86.6%. Despite a slight drop, the model retains a robust general vision reasoning ability.
@@ V Conclusion
This study seeks to reconcile the disparity between the capabilities of existing Multimodal Large Language Models (MLLMs) and the demands of robotic systems. We present ManipVQA, a novel approach designed to equip MLLMs with manipulation-centric knowledge via a visual question-answering paradigm. Our approach involves the collection of a diverse set of images featuring interactive objects, thus encompassing a broad range of challenges related to object detection, affordance, and physical concept prediction. Empirical assessments performed in robotic simulators and across various vision task benchmarks substantiate the efficacy and resilience of ManipVQA.

### v2

@@ III-C Task Formulation
III-D MLLM Fine-Tuning Strategy
III-D1 Model Architecture
We adopt the MLLM, SPHINX [13] as our primary architecture, with LLaMA2 [12] serving as the language backbone. Given the necessity for both global and local visual grounding in robotic tasks, we integrate the visual encoder from CLIP [17] to extract local semantic features and the Q-Former [31] for summarizing visual features. Spatial alignment is facilitated using projection layers, and global features are merged with local ones through channel-wise concatenation. We acknowledge that the standard image resolution for pre-trained visual encoders, typically $224\times 224$, is insufficient for detailed visual perception. This limitation is particularly significant for robotic affordance reasoning, which often requires fine-grained visual grounding of object parts, such as tool handles or machine buttons. To augment ManipVQA’s region-level grounding capabilities, we employ a sub-images patching strategy following  [13]. Specifically, we partition a $448\times 448$ image into four $224\times 224$ sub-images taken from each corner, thereby preserving intricate visual details. The resulting image tokens are then positioned before the language instructions to provide visual context for the ensuing prompts.
III-D2 Fine-Tuning Strategy
As elucidated in Sec. III-A, we model both affordance and physical concepts within natural language representations and training samples are formatted in line with the general VQA framework, as delineated in Sec. III-C. As a result, the training objective employs a unified cross-entropy loss, diverging from the approaches in  [6, 4]. To maintain the model’s broad visual reasoning proficiency, we amalgamate general visual reasoning exercises with tasks specific to robotics.
@@ IV-A Implementation Details
IV-A1 Training Details
We fine-tuned the ManipVQA model using the SPHINX framework [13] on eight NVIDIA A100 (80GB) GPUs. The fine-tuning was completed in a single epoch, which took approximately 6 hours. During this phase, the visual encoders were kept frozen to maintain the integrity of the pre-trained features. The pre-trained model was the SPHINX-1K, obtained from the official repository. Training was conducted with a batch size of 4 and a learning rate set to $2\times 10^{-5}$.
@@ IV-A1 Training Details
TABLE III: Evaluation results on HANDAL Dataset. Task abbreviations are as follows: “Obj.” for Complete Object Detection, “Aff.” for Robotic Affordance Detection, and “Gr” for Grounded Detection. The letter “B” denotes bounding box format, and “M” denotes mask representation format. Object abbreviations are listed in sequence: Hammer, Pliers-Fixed Joint, Pliers-Slip Joint, Pliers-Locking, Power Drill, Ratchet, Screwdriver, Wrench-Adjustable, Wrench-Combinational, Label, Measuring Cup, Mug, Pan, Spatula, Strainer, Utensil, and Whisk.
IV-A2 Connected with Robotic Policy
The objective of ManipVQA is to augment the generalizability of robotic control policies. It could be used in a language-free format during high-level decision-making for physically grounded knowledge. For affordance localization, while the initial bounding box identifies areas of potential manipulation, it may also encompass extraneous elements like background features. To achieve a more precise affordance map, we leverage the SAM-HQ variant [32], which uses the initial bounding box as a “box prompt” for more accurate segmentation. Heuristic methods are then employed to determine the contact point. In our experiments, the ground truth (GT) surface normals, essential for rotation estimation, were presumed to be accessible either through RGB-D sensing or an alternative pre-trained model.
@@ IV-B Experimental Setup
IV-B1 Robotic Affordance Detection
Our evaluation is primarily conducted on the HANDAL Dataset [10], assessing both bounding box average precision (AP) and pixel-wise segmentation AP. Unlike the baseline model in the HANDAL, which detects only the whole object, our ManipVQA is capable of identifying both the entire object and its manipulable parts, namely the affordance regions. Additionally, we compare our approach to LISA [33], which integrates LLM and SAM decoder.
IV-B2 Physical Concept Grounding
We utilize the PhysObjects Dataset [11] to evaluate physical concept grounding capabilities. We benchmark ManipVQA against PG-InstructBLIP [11], a fine-tuned version of InstructBLIP on PhysObjects, and the latest and most advanced MLLM, GPT-4V. Due to the limited localization capacity of GPT-4V, we pair it with Set-of-Mark [34], where bounding boxes and index numbers are explicitly annotated on the target objects within the input images.
IV-B3 General Affordance Grounding
Although ManipVQA is trained solely on a robotic affordance dataset, we are interested in its generalization capabilities on broader affordance grounding datasets, such as on AGD20K [27]. This exploration is motivated by the robust reasoning and generalization potential of LLMs. Our method is evaluated on AGD20K and follows its metrics, including KLD, SIM, and NSS. We compared with AffordanceLLM [6], Cross-View-AG [27], LOCATE [35] and 3DOI [36].
IV-B4 Robotic Manipulation Tasks
We further integrate ManipVQA with a basic robotic control policy, as detailed in Section IV-A, and use the manipulation success rate as a metric to gauge its practicality in robotic manipulation tasks using the PartNet-Mobility Dataset within the SAPIEN [37] simulator. Our experimental setup is modeled after ManipLLM [4], utilizing identical metrics. However, we employ ground truth (GT) surface normals from the simulator for rotation estimation. We compare our method with Where2Act [38], FlowBot3D [39], and ManipLLM [4]. For additional details on the experimental settings, please refer to ManipLLM [4].
@@ IV-C Results
IV-C1 Robotic Affordance Detection Evaluation
As illustrated in Table III, our ManipVQA achieves remarkable performance in the detection of both complete objects and their affordances in a unified framework. Enhanced by the SAM-HQ, our model attains superior results in tasks involving affordance detection and grounding with a mask. However, it falls short in complete object segmentation, which we attribute to a tendency of SAM-HQ to cause over-segmentation.
Figure 3: Illustration of the affordance detection and grounding tasks using the HANDAL Dataset. The first four columns show the results for affordance detection, and the last two columns present the results for the affordance grounding task. In the affordance grounding task, a description of the tool’s intended use is given instead of explicit part names. Blue masks are produced by SAM-HQ based on ManipVQA’s predicted bounding boxes, and red masks are LISA’s output.
IV-C2 Physical Concept Grounding Evaluation
Table IV presents the evaluation results on the PhysObjects [11]. It is noteworthy that even the advanced MLLM GPT-4V encounters challenges with tasks that require an understanding of physical concepts. This struggle probably stems from its limited capacity for precise localization and a deficiency in visual physical reasoning. With the integration of SoM [34], the GPT-4V has better localization ability while its performance remains suboptimal. Our ManipVQA outperforms PG-InstructCLIP [11] which is also fine-tuned on PhysObjects, and we hypothesize that this enhanced performance can be attributed to the more powerful LLM and the ensemble of vision encoders deployed within our MLLM.
@@ IV-C2 Physical Concept Grounding Evaluation
IV-C3 General Affordance Grounding
Table V presents the evaluation results on the AGD20K Dataset [27] using the Hard split as defined by [6]. Remarkably, although our ManipVQA framework is trained solely on robotic affordances, such as grasping, it still demonstrates competitive performance on the general affordance grounding task. Notably, our method achieves the best NSS score, indicating strong model performance, but also the highest KLD, which is undesirable. We attribute this high KLD to the tendency of the SAM-series models to over-segment images. Additionally, the discrepancy between the ground truth represented as a heatmap, and our output, a segmentation mask, likely contributes to the lower KLD score. Our method also shows a promising ability to distinguish between fine-grained affordances associated with the same object class. For example, it can differentiate between Cut-with and Hold actions for a knife, which involve the blade and handle, respectively, as illustrated in Fig. 5.
Figure 4: Ilustration for the general affordance grounding tasks on the AGD20K [27]. Notably, the action Sit-On is unseen in our affordance training split. Blue masks are produced by SAM-HQ based on ManipVQA’s predicted bounding boxes.
Method SIM $\uparrow$ NSS $\uparrow$ KLD $\downarrow$
@@ IV-C3 General Affordance Grounding
TABLE V: General affordance evaluation results on AGD20K [27].
Figure 5: Strong reasoning capability of ManipVQA to distinguish subtle features within affordances of the same object class. Blue masks are produced by SAM-HQ based on ManipVQA’s predicted bounding boxes.
IV-C4 Robotic Manipulation in Simulator
Table VI illustrates the zero-shot performance of our model in SAPIEN [37] when combined with a basic heuristic-based control policy. The model’s success is largely due to the retention of commonsense reasoning capabilities in the MLLM and the incorporation of affordance knowledge, which enables effective robotic manipulation without prior fine-tuning on the task-specific data.
Figure 6: Visualizatios of ManipVQA’s predictions within the SAPIEN simulator. The first sub-image depicts the experimental setup, where the model predicts the gripper’s contact point, represented by a red star, to achieve the desired movement of a specific object part. The green bounding boxes denote ManipVQA’s predictions, while the blue masks are obtained using SAM-HQ. The contact point is heuristically determined as the geometric center of the blue mask.
@@ IV-C4 Robotic Manipulation in Simulator
TABLE VI: Performance evaluation within the SAPIEN simulator using PartNet-Mobility Dataset. Notably, while the baseline methods use distinct training and testing splits, our model achieves robust performance without fine-tuning on the SAPIEN samples.
@@ IV-D Further Analysis
IV-D1 Ablation Studies
To dissect the contributions of the designs in ManipVQA, we performed a series of ablation studies. The results are presented in Table VII. It was observed that the SOTA MLLM, SPHINX, cannot execute vision-based reasoning tasks in robotics without the ManipVQA Dataset. Furthermore, without the mixture of general vision data during the fine-tuning process, there is a noticeable decline in its physical understanding and affordance reasoning capabilities. This suggests that the current ManipVQA Dataset alone may not provide sufficiently large or diverse samples for effectively fine-tuning an MLLM. Additionally, the absence of visual ensembles leads to a significant drop in the model’s ability to reason about affordances, likely because robotic affordance reasoning often requires detailed part-level understanding.
Manip. Data Vis. Ens Mix. Train Phys. Aff. Box Aff. Mask
@@ IV-D1 Ablation Studies
TABLE VII: Ablation Studies. “Manip. Data” indicates the use of the ManipVQA Dataset, “Vis. Ens” represents the employment of a visual encoder ensemble, and “Mix. Train” refers to the inclusion of a mixed general visual dataset during fine-tuning. “Phys.” assesses the model’s physical grounding; and “Aff.” denotes the model’s affordance reasoning capabilities.
IV-D2 Impact on Pre-Existing Vision Reasoning Ability
When fine-tuning a model on a specialized dataset like ManipVQA, it is essential to consider the potential impact on its pre-established general vision reasoning skills. To investigate this, we evaluated the model on the val split of RefCOCO+ [28]. The post-fine-tuning accuracy was recorded at 81.8%. This performance is to be compared against the accuracy of the pre-trained model from [2] at 86.6%. Despite a slight drop, the model retains a robust general vision reasoning ability.
@@ V Conclusion
This study seeks to reconcile the disparity between the capabilities of existing MLLMs and the demands of robotic systems. We present ManipVQA, a novel approach designed to equip MLLMs with manipulation-centric knowledge via a visual question-answering paradigm. Our approach involves the collection of a diverse set of images featuring interactive objects, thus encompassing a broad range of challenges related to object detection, affordance, and physical concept prediction. Empirical assessments performed in robotic simulators and across various vision task benchmarks substantiate the efficacy and resilience of ManipVQA.